# A YTD correspondence for constant scalar curvature metrics

### v1

@@ 5 Continuity of the radial $\beta$-entropy
From [10, Lemma A.2] it follows that
@@
Using Lemma 3.12, we can write
$d_{1}(0,u_{t}^{\beta})\leq d_{1}(0,u_{t})=-I(u_{t})\leq-(1-t)I(u_{0})-tI(u_{1}),$
where we used the simple fact that $t\mapsto I(u_{t})$ is concave. Thus,
$d_{1}(0,u_{t}^{\beta})\leq\max\{d_{1}(0,u_{0}),d_{1}(0,u_{1})\}.$
As a result, we obtain that
@@
Recall from [42, Lemma 3.2] that for $\{v_{t}\}\in\mathcal{R}^{1}$ there exists $c_{v}\in\mathbb{R}$ with $\sup_{X}v_{t}=c_{v}t$ for $t>0$, and $\{v_{t}\}\mapsto c_{v}$ is $d_{1}^{c}$-continuous. Thus, after possibly replacing $u_{t}$ and $u^{m}_{t}$ by $u_{t}-ct$ and $u^{m}_{t}-ct$ for sufficiently large $c>0$, we may assume that $u_{t},u^{m}_{t}\leq 0$.
@@
Separating the last integral, we estimate the first term. Using [10, Lemma A.2] the same way as in (21) we obtain
@@
Similarly, $\int_{X}(\max(u,v)-v)\omega^{n}_{\frac{u+v}{4}}\leq Cd_{1}(\max(u,v),v)^{\frac{1}{2^{n}}}\max(d_{1}(0,u),d_{1}(0,v))^{1-\frac{1}{2^{n}}}.$ After adding the last two estimates, the result follows from [32, Remark 5.6]. ∎
As pointed out in §1, combining Theorem 5.2 with Theorem 4.4, [26, 27] and [61], the proof of our YTD correspondence (Theorem 1.1) is now complete.
6 A transcendental YTD correspondence
@@
A Kähler test configuration $\mathcal{T}$ is a pair $\mathcal{T}:=(\mathcal{X},\mathcal{A})$ such that:
@@
$\bullet$ There exists a $\mathbb{C}^{*}$-equivariant biholomorphism
$\mathcal{X}\setminus\pi^{-1}(0)\ \cong\ X\times(\mathbb{P}^{1}\setminus\{0\}),$
identifying $\pi$ with projection onto $\mathbb{P}^{1}$.
$\bullet$ $\mathcal{A}\in H^{1,1}_{BC}(\mathcal{X},\mathbb{R})$ is a $\mathbb{C}^{*}$-invariant Bott-Chern cohomology class, and
$\mathcal{A}|_{X\times(\mathbb{P}^{1}\setminus\{0\})}=p_{1}^{*}[\omega],$
where $p_{1}$ is projection to $X$.
@@
As in the ample case, it is possible to associate to $\mathcal{T}$ a $C^{1,1}$ geodesic ray $\{u^{\mathcal{T}}_{t}\}$ ([79, Lemma 4.6], [30]), allowing to define $K^{\beta}$-stability in this context in the same way as in the ample case (recall (9)).
A combination of [65, Proposition 6.3.1] and [65, Theorem 5.1.7] implies that finite energy geodesic rays with bounded K energy slope are $d_{1}^{c}$-approximable by rays of Kähler test configurations. As a result, the exact same proof as that of Theorem 1.1 implies the following:
@@ 7 A slope formula for the $\beta$-entropy
where the sup is over all geodesic rays $\{v_{t}\}$ arising from semi-ample test-configurations of $(X,L)$, and
@@
The boundedness assumption on $\{u_{t}\}$ is naturally satisfied when it comes to applications, since it holds for a large class of geodesic rays (including the ones induced from filtrations; see [76, §9]). That being said, it would be nice to know if this condition is necessary.
@@
Since we are dealing with sublinear subgeodesic rays, we can assume that $t\to u_{t}$ and $t\to v_{t}$ are both $t$-decreasing. Also, [33, Theorem 1] gives some $\gamma>0$ such that
@@
For this step we fix a geodesic ray $\{u_{t}\}\in\mathcal{R}^{1}$ normalized by $\sup u_{t}=0$ (this is possible due to [40, Lemma 3.2]).
@@
where in the second inequality we used the convexity of $t\mapsto d_{1}(\phi_{j,t,\beta},u_{t})$ [11, Proposition 5.1] and in the second line we used the linearity of $t\mapsto I(\phi_{j,t,\beta})-I(u_{t})$.
@@
Due to (32) we have that $I(\phi_{t,\beta})-I(u_{t})=I(\phi_{t,\beta})+I(u_{t})-2I(P(u_{t},\phi_{t,\beta}))$. Using the $\mathcal{E}^{1}$ version of [33, Proposition 2.3] we obtain that $P(u_{t},\phi_{t,\beta})=u_{t}$. From this and the fact that $u^{\beta}_{t_{j}}\leq 0$ we obtain that $\phi_{t,\beta}\leq 0$ and
@@
A byproduct of the above construction is the following important theorem. In the setting of algebraic quantization, analogous results have been conjectured by the second author in [95, Conjecture 3.3], but are still open to our knowledge.
Theorem 7.3 (=Theorem 1.2).
Let $\{u_{t}\}\in\mathcal{R}^{1}$. Then
$\lim_{\beta\to\infty}K^{\beta}\{u_{t}\}=K\{u_{t}\}.$
@@
Now sending $\beta\to\infty$ and using that $K\{\cdot\}$ is $d_{1}^{c}$-lsc ([25, Proposition 5.9]) we obtain that $\lim_{\beta\to\infty}K^{\beta}\{u_{t}\}=K\{u_{t}\}.$
If $Ent\{u_{t}\}=\infty$, we claim that $\lim_{\beta\to\infty}Ent^{\beta}\{u_{t}\}=\infty$ as well. Indeed, suppose otherwise, then Proposition 3.10 implies that $Ent^{\beta}\{u_{t}\}\leq C$ for some $C>0$ independent of $\beta>0$. So (33) implies that $d_{1}^{c}(\{\phi_{t,\beta}\},\{u_{t}\})\to 0$ as $\beta\to\infty$. Sending $\beta\to\infty$ in (34) and using that $K\{\cdot\}$ is $d_{1}^{c}$-lsc and $K^{\beta}\{u_{t}\}\leq C$, we obtain $K\{u_{t}\}<\infty$, a contradiction. So we conclude. ∎
@@
be the smallest geodesic ray lying above $\{u_{t,\beta}\}$. Such $\{v_{t,\beta}\}$ can be explicitly constructed as the increasing limit of the geodesic segments $0,k]\ni t\mapsto\phi_{k}(t)$ joining 0 and $u_{k,\beta}$ (for details, see the proof of [42, Proposition 3.13]).
@@
Then by Corollary 4.5 and (the proof of) (26),
$K\{v_{t,\beta}\}<\infty.$
Thus, $\{v_{t,\beta}\}$ must be maximal in the sense of [10], due to [61, Theorem 1.2]. Since $\{v_{t,\beta}\}$ is the smallest possible ray dominating the subgeodesic $\{u_{t,\beta}\}$, in the jargon of [10], $\{v_{t,\beta}\}$ is the maximal geodesic ray associated to the subgeodesic $\{u_{t,\beta}\}$.
Thus the following estimate holds, thanks to [61, Lemma 3.1]:
@@
[61, Lemma 3.1] proves this for geodesic rays, but the same argument works for the subgeodesic ray $\{u_{t,\beta}\}$ (c.f. (15)), since $u_{t,\beta}\leq 0$ due to Lemma 3.1(ii) and $u_{t}\leq 0$ . ∎
@@
So we conclude by the comparison principle [6, Lemma 2.5]. ∎
Let us fix $\delta\in(0,1)$:
$\delta:=\frac{A}{A+\beta}.$
Then $\frac{\beta}{1-\delta}=\beta+A$ and Proposition 3.14 implies that $t\mapsto((1-\delta)u_{t})^{\frac{\beta}{1-\delta}}=((1-\frac{A}{\beta+A})u_{t})^{\beta+A}$ is a subgeodesic ray for any $\beta>0$. By our notation (36), the smallest geodesic ray lying above it was denoted by $\{v_{t,\beta+A}\}$.
We claim that $(1-\delta)\phi_{t,\beta}\leq v_{t,\beta+A}.$ Indeed, recall that $\phi_{t,\beta}$ is constructed as the limit of geodesic segments $\phi_{j,t,\beta}$ joining 0 and $u^{\beta}_{t_{j}}$. Then
$0,t_{j}]\ni t\mapsto(1-\delta)\phi_{j,t,\beta}$
is a subgeodesic segment joining 0 and $(1-\delta)u_{t_{j}}^{\beta}$. Due to Lemma 7.7 and the comparison principle of weak geodesics, this subgeodesic is bounded from above by the geodesic segment joining 0 and $((1-\delta)u_{t_{j}})^{\frac{\beta}{1-\delta}}-\frac{n(1-\delta)}{\beta}\log(1-\delta)$. Sending $j\to\infty$, the constant term $\frac{n(1-\delta)}{\beta}\log(1-\delta)$ will disappear in this limit process, and we obtain the claimed key relation:
@@
For any $\beta>1$ and any bounded geodesic ray $\{u_{t}\}$, there exists a sequence of geodesic rays $\{\psi_{t,k}\}$ such that
$Ent^{\beta}\{u_{t}\}=\lim_{k\to\infty}(L^{\beta}(\{\psi_{t,k}\},\{u_{t}\})+\beta(I\{\psi_{t,k}\}-I\{u_{t}\})).$
Moreover, the sequence $\{\psi_{t,k}\}$ can be chosen so that $Ent\{\psi_{t,k}\}<\infty,$
@@
$\beta(I\{\phi_{t,\beta_{\varepsilon}}\}-I\{u_{t}\})=(1-\varepsilon)^{2}Ent^{\frac{\beta}{(1-\varepsilon)^{2}}}\{u_{t}\}\geq(1-\varepsilon)^{2}Ent^{\beta}\{u_{t}\}.$ (39)
Moreover, by Corollary 3.7, $Ent^{\frac{\beta}{(1-\varepsilon)^{2}}}\{u_{t}\}\leq\frac{1}{(1-\varepsilon)^{2}}Ent^{\beta}\{u_{t}\}.$ Theorefore,
@@
for some $\beta>0$ and $\varepsilon>0$. Then there exists a sequence of geodesic rays $\{w_{t,m}\}$ induced from semi-ample test configurations such that
$I\{w_{t,m}\}\to I\{\psi_{t}\}\text{ as }m\to\infty,$
and
$\int_{0}^{\infty}e^{-4\varepsilon t}\int_{X}e^{\beta(w_{t,m}-u_{t})}\omega^{n}dt<\infty\text{ for all }m\gg 1.$
@@
During the proof, $\beta>0$ and $\varepsilon>0$ will be fixed. Since $Ent\{\psi_{t}\}<\infty$, [61, Theorem 1.2] is applicable, and applying Demailly’s regularization to $\{\psi_{t}\}$ gives a sequence of geodesic rays $\{w_{t,m}\}$ induced from semi-ample test configurations such that $d_{1}^{c}(\{w_{t,m}\},\{\psi_{t}\})\to 0$, hence
@@
for all $m\gg 1$. We can and will assume that
$m_{0}>\beta.$
@@
After possibly adding a linear term, we can assume that $\sup u_{t}=0$ [40, Lemma 3.2]. Due to the definition of $Ent^{\beta}$ (18) it suffices to find a sequence of geodesic rays $\{w_{t,j}\}$ arising from semi-ample test-configurations such that
@@
Additionally, by Proposition 7.9, for each $k>0$, we can find $m_{k}\gg 1$ and a geodesic ray $\{w_{t,m_{k}}\}$ arising from a semi-ample test configuration such that
@@
In this section, the polarization assumption for $(X,\omega)$ is only used to invoke Demailly’s approximation when proving Lemma 7.5 and Proposition 7.9. However, as explained in §6, such an approximation procedure also holds for the general Kähler case, thanks to [65]. As a consequence, the transcendental version of Theorem 7.1 holds as well, where the sup in (28) is then over all geodesic rays $\{v_{t}\}$ attached to Kähler test configurations.
8 Algebraic interpretation of $K^{\beta}$-stability
Assume that $(X,L)$ is a polarized manifold. Let $\omega\in c_{1}(L)$, so $V=L^{n}$. Assume that $\{u_{t}\}$ is a geodesic ray arising from a semi-ample test configuration $\mathcal{T}=(\mathcal{X},\mathcal{L})$. Then we write

### v2

@@ 5 Continuity of the radial $\beta$-entropy
From [9, Lemma A.2] it follows that
@@
Using Lemma 3.12, we can write $d_{1}(0,u_{t}^{\beta})\leq d_{1}(0,u_{t})=-I(u_{t})\leq-(1-t)I(u_{0})-tI(u_{1}),$ where we used the simple fact that $t\mapsto I(u_{t})$ is concave. Thus, $d_{1}(0,u_{t}^{\beta})\leq\max\{d_{1}(0,u_{0}),d_{1}(0,u_{1})\}.$ As a result, we obtain that
@@
Recall from [46, Lemma 3.2] that for $\{v_{t}\}\in\mathcal{R}^{1}$ there exists $c_{v}\in\mathbb{R}$ with $\sup_{X}v_{t}=c_{v}t$ for $t>0$, and $\{v_{t}\}\mapsto c_{v}$ is $d_{1}^{c}$-continuous. Thus, after possibly replacing $u_{t}$ and $u^{m}_{t}$ by $u_{t}-ct$ and $u^{m}_{t}-ct$ for sufficiently large $c>0$, we may assume that $u_{t},u^{m}_{t}\leq 0$.
@@
Separating the last integral, we estimate the first term. Using [9, Lemma A.2] the same way as in (35) we obtain
@@
Similarly, $\int_{X}(\max(u,v)-v)\omega^{n}_{\frac{u+v}{4}}\leq Cd_{1}(\max(u,v),v)^{\frac{1}{2^{n}}}\max(d_{1}(0,u),d_{1}(0,v))^{1-\frac{1}{2^{n}}}.$ After adding the last two estimates, the result follows from [36, Remark 5.6]. ∎
In conclusion, all the ingredients necessary for the proof of Theorem 1.1, as presented in the introduction, have now been established.
A transcendental YTD correspondence
@@
A Kähler test configuration $\mathcal{T}$ of $(X,\omega)$ is a pair $\mathcal{T}:=(\mathcal{X},\mathcal{A})$ such that:
@@
$\bullet$ There exists a $\mathbb{C}^{*}$-equivariant biholomorphism $\mathcal{X}\setminus\pi^{-1}(0)\ \cong\ X\times(\mathbb{P}^{1}\setminus\{0\}),$ identifying $\pi$ with projection onto $\mathbb{P}^{1}$.
$\bullet$ $\mathcal{A}\in H^{1,1}_{BC}(\mathcal{X},\mathbb{R})$ is a $\mathbb{C}^{*}$-invariant Bott-Chern cohomology class, and $\mathcal{A}|_{X\times(\mathbb{P}^{1}\setminus\{0\})}=p_{1}^{*}[\omega],$ where $p_{1}$ is projection to $X$.
@@
As in the ample case, it is possible to associate to $\mathcal{T}$ a $C^{1,1}$ geodesic ray $\{u^{\mathcal{T}}_{t}\}$ ([90, Lemma 4.6], [34]), allowing to define $K^{\beta}$-stability in this context in the same way as in the algebraic case (recall (10)).
A combination of [75, Proposition 6.3.1] and [75, Theorem 5.1.7] implies that finite energy geodesic rays with bounded $K$ energy slope are $d_{1}^{c}$-approximable by rays of Kähler test configurations. As a result, the exact same proof as that of Theorem 1.1 implies the following:
@@ 6 A slope formula for the $\beta$-entropy
where the sup is over all geodesic rays $\{v_{t}\}$ arising from test-configurations of $(X,L)$, and
@@
The boundedness assumption on $\{u_{t}\}$ is naturally satisfied when it comes to applications, since it holds for a large class of geodesic rays (including the ones induced from filtrations; see [86, §9]). That being said, it would be nice to know if this condition is necessary.
@@
Since we are dealing with sublinear subgeodesic rays, we can assume that $t\to u_{t}$ and $t\to v_{t}$ are both $t$-decreasing. Also, [37, Theorem 1] gives some $\gamma>0$ such that
@@
For this step we fix a geodesic ray $\{u_{t}\}\in\mathcal{R}^{1}$ normalized by $\sup u_{t}=0$ (this is possible due to [44, Lemma 3.2]).
@@
where in the second inequality we used the convexity of $t\mapsto d_{1}(\phi_{j,t,\beta},u_{t})$ [10, Proposition 5.1] and in the second line we used the linearity of $t\mapsto I(\phi_{j,t,\beta})-I(u_{t})$.
@@
Due to (46) we have that $I(\phi_{t,\beta})-I(u_{t})=I(\phi_{t,\beta})+I(u_{t})-2I(P(u_{t},\phi_{t,\beta}))$. Using the $\mathcal{E}^{1}$ version of [37, Proposition 2.3] we obtain that $P(u_{t},\phi_{t,\beta})=u_{t}$. From this and the fact that $u^{\beta}_{t_{j}}\leq 0$ we obtain that
@@
A byproduct of the above construction is the following important theorem. In the setting of algebraic quantization, analogous results have been conjectured by the second author in [106, Conjecture 3.3], but are still open to our knowledge.
Theorem 6.3.
Let $\{u_{t}\}\in\mathcal{R}^{1}$. Then $\lim_{\beta\to\infty}K^{\beta}\{u_{t}\}=K\{u_{t}\}.$
@@
Now sending $\beta\to\infty$ and using that $K\{\cdot\}$ is $d_{1}^{c}$-lsc ([29, Proposition 5.9]) we obtain that $\lim_{\beta\to\infty}K^{\beta}\{u_{t}\}=K\{u_{t}\}.$
If $Ent\{u_{t}\}=\infty$, we claim that $\lim_{\beta\to\infty}Ent^{\beta}\{u_{t}\}=\infty$. Indeed, suppose otherwise, then Proposition 3.10 implies that $Ent^{\beta}\{u_{t}\}\leq C$ for some $C>0$ independent of $\beta>0$. So (47) implies that $d_{1}^{c}(\{\phi_{t,\beta}\},\{u_{t}\})\to 0$ as $\beta\to\infty$. Sending $\beta\to\infty$ in (48) and using that $K\{\cdot\}$ is $d_{1}^{c}$-lsc and $K^{\beta}\{u_{t}\}\leq C$, we obtain $K\{u_{t}\}<\infty$, a contradiction. So we conclude. ∎
@@
be the smallest geodesic ray lying above $\{u_{t,\beta}\}$. Such $\{v_{t,\beta}\}$ can be explicitly constructed as the increasing limit of the geodesic segments $0,k]\ni t\mapsto\phi_{k}(t)$ joining 0 and $u_{k,\beta}$ (for details, see the proof of [46, Proposition 3.13]).
@@
Then by Corollary 4.5 and (the proof of) (40), $K\{v_{t,\beta}\}<\infty.$ Thus, $\{v_{t,\beta}\}$ must be maximal in the sense of [9], due to [70, Theorem 1.2]. Since $\{v_{t,\beta}\}$ is the smallest possible ray dominating the subgeodesic $\{u_{t,\beta}\}$, in the jargon of [9], $\{v_{t,\beta}\}$ is the maximal geodesic ray associated to the subgeodesic $\{u_{t,\beta}\}$.
Thus the following estimate holds, thanks to [70, Lemma 3.1]:
@@
[70, Lemma 3.1] proves this for geodesic rays, but the same argument works for the subgeodesic ray $\{u_{t,\beta}\}$ (c.f. (22)), since $u_{t,\beta}\leq 0$ due to Lemma 3.1(ii) and $u_{t}\leq 0$ . ∎
@@
So we conclude by the comparison principle [5, Lemma 2.5]. ∎
Let us fix $\delta\in(0,1)$ by $\delta:=\frac{A}{A+\beta}.$ Then $\frac{\beta}{1-\delta}=\beta+A$ and Proposition 3.14 implies that $t\mapsto((1-\delta)u_{t})^{\frac{\beta}{1-\delta}}=((1-\frac{A}{\beta+A})u_{t})^{\beta+A}$ is a subgeodesic ray for any $\beta>0$. By our notation (50), the smallest geodesic ray lying above it was denoted by $\{v_{t,\beta+A}\}$.
We claim that $(1-\delta)\phi_{t,\beta}\leq v_{t,\beta+A}.$ Indeed, recall that $\phi_{t,\beta}$ is constructed as the limit of geodesic segments $\phi_{j,t,\beta}$ joining 0 and $u^{\beta}_{t_{j}}$. Then $0,t_{j}]\ni t\mapsto(1-\delta)\phi_{j,t,\beta}$ is a subgeodesic segment joining 0 and $(1-\delta)u_{t_{j}}^{\beta}$. Due to Lemma 6.7 and the comparison principle of weak geodesics, this subgeodesic is bounded from above by the geodesic segment joining 0 and $((1-\delta)u_{t_{j}})^{\frac{\beta}{1-\delta}}-\frac{n(1-\delta)}{\beta}\log(1-\delta)$. Sending $j\to\infty$, the constant term $\frac{n(1-\delta)}{\beta}\log(1-\delta)$ will disappear in this limit process, and we obtain the claimed key relation:
@@
For any $\beta>1$ and any bounded geodesic ray $\{u_{t}\}$, there exists a sequence of geodesic rays $\{\psi_{t,k}\}$ such that $Ent^{\beta}\{u_{t}\}=\lim_{k\to\infty}(L^{\beta}(\{\psi_{t,k}\},\{u_{t}\})+\beta(I\{\psi_{t,k}\}-I\{u_{t}\})).$ Moreover, the sequence $\{\psi_{t,k}\}$ can be chosen so that $Ent\{\psi_{t,k}\}<\infty,$
@@
$\beta(I\{\phi_{t,\beta_{\varepsilon}}\}-I\{u_{t}\})=(1-\varepsilon)^{2}Ent^{\beta_{\varepsilon}}\{u_{t}\}\geq(1-\varepsilon)^{2}Ent^{\beta}\{u_{t}\}.$ (53)
Moreover, by Corollary 3.7, $Ent^{\beta_{\varepsilon}}\{u_{t}\}\leq\frac{1}{(1-\varepsilon)^{2}}Ent^{\beta}\{u_{t}\}.$ Theorefore,
@@
for some $\beta>0$ and $\varepsilon>0$. Then there exists a sequence of geodesic rays $\{w_{t,m}\}$ induced from test configurations such that
$\lim_{m\to\infty}I\{w_{t,m}\}=I\{\psi_{t}\}\text{ and }\int_{0}^{\infty}e^{-4\varepsilon t}\int_{X}e^{\beta(w_{t,m}-u_{t})}\omega^{n}dt<\infty\text{ for all }m\gg 1.$
@@
During the proof, $\beta>0$ and $\varepsilon>0$ will be fixed. Since $Ent\{\psi_{t}\}<\infty$, [70, Theorem 1.2] is applicable, and applying Demailly’s regularization to $\{\psi_{t}\}$ gives a sequence of geodesic rays $\{w_{t,m}\}$ induced from test configurations such that $d_{1}^{c}(\{w_{t,m}\},\{\psi_{t}\})\to 0$, hence
@@
for all $m\gg 1$. We can and will assume that $m_{0}>\beta.$
@@
After possibly adding a linear term, we can assume that $\sup u_{t}=0$ [44, Lemma 3.2]. Due to the definition of $Ent^{\beta}$ (32) it suffices to find a sequence of geodesic rays $\{w_{t,j}\}$ arising from test-configurations such that
@@
Additionally, by Proposition 6.9, for each $k>0$, we can find $m_{k}\gg 1$ and a geodesic ray $\{w_{t,m_{k}}\}$ arising from a test configuration such that
@@
In this section, the polarization assumption for $(X,\omega)$ is only used to invoke Demailly’s approximation when proving Lemma 6.5 and Proposition 6.9. However, such an approximation procedure also holds for the general Kähler case, thanks to [75, Proposition 6.3.1]. As a consequence, the transcendental version of Theorem 6.1 holds as well, where the sup in (42) is then over all geodesic rays $\{v_{t}\}$ attached to Kähler test configurations.
7 Threshold type interpretation of $K^{\beta}$-stability
Assume that $(X,L)$ is a polarized manifold. Let $\omega\in c_{1}(L)$, so $V=L^{n}$. Assume that $\{u_{t}\}$ is a geodesic ray arising from a test configuration $\mathcal{T}=(\mathcal{X},\mathcal{L})$. Then we write